OpenStudy (anonymous):

PB has endpoints at –2 and 12. What is the coordinate of its midpoint?

OpenStudy (anonymous):

(-2+12)/2=10/2 =5

OpenStudy (anonymous):

mandolino is correct it is 5

OpenStudy (anonymous):

:) thanks